# Towards Foundation Models: Evaluation of Geoscience Artificial Intelligence with Uncertainty

### v1

@@ 3 Methods - Model Architecture, 3 Training Approaches, and 5 Models
We train all models using a sufficient number of epochs until convergence. In general, the Standard models reached convergence within 50 epochs and the TL models reached convergence within 12 epochs. We selected the model instance for evaluation that exhibited lowest loss on the validation data within a cosine annealing cycle.
@@ 4.1 Metrics
As our main goal is to demonstrate the effectiveness and versatility of our evaluation framework, we limit the scope of evaluation of the phase picking models by considering just a few classification metrics and one functional based regression metric. To form these metrics, we first implement techniques similar to those found in [47, 30, 20] to obtain model predictions for test waveforms. The predictions are probabilities of the P-arrival for each sample in each waveform. From the predicted curves and through an algorithm similar to those found in [ref same papers as above], we obtain P-arrival picks, which are the discrete locations in time that a P-arrival is predicted by the model. There can be none, one, or any number of picks per waveform. To motivate the specific metrics we consider, we note that it is important in evaluation of phase pickers to consider both the model’s strength in identifying if a P-arrival occurs (e.g., alerting) and, if so, the associated accuracy of the pick (e.g., how close is the predicted pick to the true P-arrival).
Evaluating identification is typically done using classification metrics defined using the picks [47, 30, 20, 45]. True positives ($TP$) are assigned to picks if they are near enough to the labeled P-arrival. False positives ($FP$) are assigned to picks either far from the labeled P-arrival or to extra picks that are not true positives. False negatives ($FN$) occur if no picks occur near the labeled P-arrival. True negatives are not informative in this context and are not recorded. In our display, we consider summary metrics including $Recall=\frac{TP}{TP+FN}$ and $F1=\frac{2TP}{2TP+FP+FN}$ which are only applied to earthquake waveforms. An example of a model prediction featuring classified picks including a true positive and 3 false positives is shown in Figure 3. For noise waveforms, we calculate noise percent correct as the number of noise waveforms containing no predicted picks ($TN_{noise}$) divided by the total number of noise waveforms ($n_{noise}$). That is, $Noise\,\%\,Correct=\frac{TN_{noise}}{n_{noise}}$.
@@
Figure 3: The predicted probability of the P-arrival for the waveform shown in Figure 1. The labeled P-arrival is displayed on the x-axis as a blue triangle along with 4 predicted picks (green stars) along with labels that classify the pick as either a false positive or true positive. An inset plot zooms in on the labeled pick to show the true positive and an extra nearby pick classified as a false positive.
@@ 4.2 Statistical Framework for Evaluation
In this section, we present a statistical framework for comparing models in terms of their performance and uncertainty arising from data sampling and model training. The statistical framework can be thought of as a mixed effect model or a completely randomized design with sub-sampling. It allows us to distinguish performance effects with uncertainty between the 5 models and different training amounts. Furthermore, we can segregate and estimate the two sources of variation with uncertainty. We describe our overall approach here and relegate many details of the calculations to the appendix.
@@
where $m=\{1\ldots 5\}$ indexes the 5 models, $a=\{1,\ldots,5\}$ indexes the 5 quantities of clusters (representing the amount of data) at levels $\{1,3,6,9,12\}$, $d=\{1,\ldots,12\}$ indexes the randomly selected cluster set (i.e., the data choice selected to train the model), and $i=\{1,2,3,4\}$ indexes the 4 model initializations.
In this statistical model, $\gamma$ is the grand mean, $\mu_{m}$ is the effect of model $m$, $\alpha_{a}$ is the effect of the $a$’th quantity of clusters, and $\theta_{ma}$ is the interaction term. The uncertainty associated with data selection and model training is segregated into two independent random terms each assuming a zero mean independent and identically distributed Gaussian with variance inherent to a given model and quantity of clusters: $\epsilon^{data}_{mad}\sim\mathcal{N}(0,\sigma^{2data}_{ma})$ and $\epsilon^{train}_{madi}\sim\mathcal{N}(0,\sigma^{2train}_{ma})$. This is reasonable because we desire to understand how the uncertainty is impacted by the model choice and quantities of clusters used.
Using these models, we estimate the effects on the metric for each model and quantity of clusters by computing the mean over the 12 cluster sets and the 4 initializations, denoted as $\bar{y}_{ma\cdot\cdot}$. Then, we compute $90\%$ confidence intervals around those estimates using the Gaussian assumptions. Note that the intervals are influenced more heavily by the cluster set (data choice) uncertainty because at each cluster set, we average over the initializations thereby reducing training uncertainty influence.
The above statistical model and resulting confidence intervals rely on Gaussian assumptions for the two error terms. That is, the distribution of the error terms implies a distribution of the metric as $y_{madi}\sim\mathcal{N}(\mu_{m}+\alpha_{a}+\theta_{ma},\sigma^{2data}_{ma}+\sigma^{2train}_{ma})$. We recognize that metrics such as recall are bounded between $[0,1]$, so the Gaussian assumption theoretically cannot hold. However, in practice we have a large number of test waveforms to form these metrics and we find that the variances of these metrics are small compared to the range of the metrics’ bounds.
@@ 5 Results
After training the 1,200 model instances, computing the metrics for evaluation, and organizing them through the statistical framework, we are ready to compare the performance of the models.
@@
The classification based summary metrics of recall and noise percent correct are presented in Figure 4 along with 90% confidence intervals for the standard and 2 masked TL models showing their performance as a function of quantity of clusters used in training. All models improve in recall as more data is added with largest gains between 1 and 3 clusters followed by diminishing returns from there on. The trend is less consistent for noise percent correct where the TL w/ Frozen Weights Masked model performs slightly worse with increasing data. The variability in performances decreases for all models as more training data is added.
@@
Figure 4: The statistical framework is used to summarize classification results including earthquake recall and noise percent correct as a function of the number of training clusters. Vertical bars represent 90% confidence intervals. Largest performance gains occur as the quantity of clusters increases from 1 to 3 and crossing lines indicate some training approaches are superior for different budgets of training data.
@@
The functional regression based metric of cumulative RMSR demonstrating the evaluation’s framework versatility and importance is shown in Figure 5 for the standard and 2 TL masked models with 90% confidence intervals (ribbons) for selected quantities of clusters of 1, 6, and 12. Stark performance differences can be seen between each model especially as more training data is added as seen through tightening confidence intervals.
@@
In both Figures 4 and 5, we see that there is no clear global winner in terms of learning efficiency. The TL w/ Free Weights Masked model appears superior for recall and cumulative RMSR for the smallest training data size of 1 cluster, but is surpassed by the standard model for greater than 3 clusters. In contrast to expectations, adding more data actually degrades the TL w/ Frozen Weights Masked model for noise percent accuracy. These results would help practitioners choose the model for their given problem. Depending on the availability of training data and how they prioritize performance metrics, their choice would change.
@@
Figure 5: The statistical framework is used to summarize the regression-based functional metrics of cumulative RMSR for 1, 6, and 12 clusters (top, middle, and bottom respectively) for the standard and 2 TL masked models. Ribbons indicate 90% confidence intervals. The relative positions of the model curves change as more training data is added indicating that no model is a globally most efficient learner.
Valuable to AI evaluation is determining whether there is an effect of overlap between training and test data. To explore this aspect, we compute statistical contrasts (differences) between the unmasked and masked models. We show these contrasts in Figure 6 computed for recall, F1, and noise percent correct for the two TL models at each data amount with 90% confidence intervals. Vertical lines are drawn at the origin and if the interval contains 00, then the effect of overlap may be considered statistically significant at the 10% level. Positive values indicate that including overlap during training improves model performance.
In analyzing Figure 6, we see minimal impact from overlap for 1 cluster, but loosely positive trends with increasing training data for recall and F1. For noise percent correct, however, there appears to be little impact of overlap regardless of the training data amount. These results demonstrate that overlap should be considered in future DLM/FM evaluation because overlapping may inflate performance metrics and fair evaluation should account for such bias.
Figure 6: Contrasts between Unmasked and Masked models with 90% confidence intervals shown for the TL w/ Free Weights (red-orange; top 3 panels) and TL w/ Frozen Weights (sky-blue; bottom 3 panels) models for recall (left), F1 (center), and noise percent correct (right) evaluated at each quantity of cluster. The figure shows modest performance gains as a result of training-test overlap.
@@
Our evaluation framework allows for estimation of two sources of performance variation including random data sampling and stochastic training processes. As an example, Figure 7 shows these estimates for F1 for the standard and 2 TL masked models across the quantities of training clusters with 90% confidence intervals. We see that training and data variance decrease as more training data is added and we see some differentiation in the variability for each model. The data variance (middle panel) shrinks quickly towards zero as more data is added and is often estimated as negative. These negative estimates are not shown because negative variance estimates are theoretically impossible. This is a known challenge in these classical statistical frameworks, but negative values essentially mean the data variance component is very small relative to the training variance.
@@
The final panel in Figure 7 shows the relative contributions of each source of variation through bar plots where the opaque and transparent bars are the contributions of data and training variances, respectively. The total variance for all models trends downward with more data with diminishing returns. The variation is mostly dominated by the stochastic training processes for quantities of clusters greater than 3. Practitioners can use this information to target specific sources of variation that they suspect might dominate in their problem, perhaps by running more initializations or labeling data more strategically.
@@
Figure 7: Training variance (left) and data sampling variance (middle) are estimated with 90% confidence intervals (vertical bars) for F1 for the standard and 2 TL masked models. The right panel shows bar plots constructed as the sum of the estimates of training and data variances. The different transparency levels show the contribution of each component where the training contribution is transparent and the data contribution is opaque. Estimating performance variation and attributing the source is informative to practitioners during model selection for setting performance expectations and risk mitigation.
@@
We provide an example that emphasizes the risk of evaluating models without considering variation in performance. We consider the scenario where a practitioner is deciding among 3 models (the standard model and the 2 TL masked models) to incorporate into their workflow. They are equipped with a single cluster set, say 1 cluster’s worth of data. To make their decision, we imagine that they train one instance of each model using the fixed data. We did exactly this in our training design, but we went further by training each model 4 times under different initializations, not just once. As such, we imagine the practitioner is equally likely to get one of the four recall scores we obtained for each model. Thus, with three models and 4 equally likely results for each, there are $644^{3}=$ 64 equally likely possible outcomes that the practitioner could see. If the practitioner ranks each model based on recall score (e.g., first place, second place, third place), due to random training variation, we might expect these rankings to change if the practitioner ran the experiment a second time. The important question then becomes, what is the probability that each model gets each rank?
@@
Using our 3 models each initialized 4 times for a given quantity of clusters, we calculate the exact probabilities that each model is ranked in first, second, or third place. Furthermore, since we have 12 cluster sets for each quantity of clusters, we conduct the calculation $12$ times and take the mean to mitigate data bias. The results of this simulation are presented in Figure 8, where the probability estimates of ranking each model based on recall in first, second, or third place are shown for $1$ and $12$ cluster quantities.
@@
Figure 8 shows that, due to training and data uncertainty, there is substantial risk of choosing a model that performs more poorly on average. For example, for 1 cluster (top panel), there is a $1.0\%$ chance that the TL w/ Frozen Weights Masked (orange) model has the best recall score. In light of the recall results in Figure 4 (left panel) for 1 cluster, we know that the TL w/ Frozen Weights Masked performs substantially worse than the other two models on average. Based purely on recall, it would be a mistake if the practitioner were to choose this model for their workflow. In a similar light, claiming that the TL w/ Frozen Weights Masked model is the state of the art compared to the other two would be an even graver mistake. Our framework guards against such mistakes.
We show the probability of rankings results for the 1 and 12 clusters to show how the probability of mis-ranking is a greater risk in the lower data regime and the risk is somewhat reduced as more data is added. From these results, we learn that researchers should apply the same principles in our framework when evaluating FMs. This is because FMs are expected to need minimal labeled training data, but teasing out exactly how much data is needed to avoid the risk of mis-ranking aids in determining which model is best.
@@
Figure 8: To demonstrate the importance of considering uncertainty in the evaluation framework, the probability of ranking the standard and 2 TL masked models based on recall in first, second, or third place is shown for 1 (top) and 12 (bottom) quantities of clusters. Numbers/bars indicate the probability of ranking the associated model in each place if one were to only train one model instance. Numbers less than 100% indicate that mis-ranking the models is a risk which is detrimental to model selection.
@@ 6 Discussion
We present results that will be increasingly relevant for developers of DL pickers and for the seismic community as they invest in FMs. In inspecting the results presented in Figure 4, we see overall poor performance in the TL w/ Frozen Weight models. We hypothesized that freezing half the model at the learned weights from the global dataset would produce the best model for small training data amounts. However, the results negate this hypothesis for datasets comprising 10,000 waveforms (approximately the number of waveforms in a single cluster) or more for the PhaseNet architecture. In a similar vein, we see that models trained from scratch perform better than TL when the data size approaches 30,000 waveforms (3 clusters). Since FMs are pre-trained using a large general dataset and fine-tuned (typically weights for the last couple of layers are set to be updated) using tailored datasets to accomplish downstream tasks, we see the importance of continuing to assess the utility of seismic FMs by comparing to traditional modeling approaches.
@@
Another valuable result for the seismic AI community is that expending effort to acquire larger and larger labeled training datasets may be wasteful. This is exhibited by Figure 4 where we see diminishing improvement as the amount of training data increases. This suggest relatively smaller datasets ($\sim$30,000 or $\sim$60,000 waveforms) may be sufficient for reasonable model performance even when training models from scratch. This is a good news because large labeled datasets are expensive and sometimes unattainable. Furthermore, these results fortify the importance of learning efficiency as an aspect of future FM evaluation. Since the best model depends on the amount of training data, knowing which model or modeling approach does more with less is valuable information during model selection. And, as we have shown, it is not a simple one-model-takes-all, but rather a dynamic choice for a given situation.
As we work toward an evaluation framework with uncertainty for geoscience FMs, we recognize the limitations in using deep ensembles as the primary form for obtaining uncertainty estimates. Deep ensembles are computationally expensive, requiring at least some retraining. With FMs, unfortunately this expense is likely prohibitively large for retraining the base model. However, this limitation does not detract from the importance of quantifying training induced variation, and alternative approaches are possible such as using ensembles during fine tuning.
We have demonstrated evidence that overlap in the training/test data can imbue performance gain, but we recognize that the effect may be confounded by the 4% decrease in the amount of data in the STEAD dataset with Italy removed. In other words, it is unclear if the performance drop in the masked models are caused by removing data from the Italian region, or if the drop is simply caused by the withholding of data from any region. To tease out these confounded effects, we could train models using random regional masking, then compare their performances to the performance of the Italy masked models. If the Italy masked models are significantly worse, then we may more confidently conclude whether the cause is the Italian region.
Regarding constructing splits to measure learning efficiency, location based clustering of the INSTANCE data was a convenient and effective first order attempt for making clusters that represent different aspects of the INSTANCE data that can be used for future frameworks. However, other methods that can characterize and group the INSTANCE data based on some inherent characteristics/features could better limit leakage between the clusters. These methods could utilize other metadata such as source to station distance, earthquake type, geological characteristics of the area, or signal-based aspects. A potential drawback of these methods is that they may result in fewer clusters, thereby reducing the number of cluster sets and making the data variance more difficult to estimate.
@@
Related to data variance and clustering is that some data variances are estimated as negative because they are small relative to the training variance. One way to more precisely estimate data variance is to increase the number of cluster sets (i.e. from 12 randomly chosen cluster sets to 30). However, we would not see much variation in the cluster sets for large quantities of clusters (e.g., 9 or 12) because we do not sample with replacement. Sampling with replacement would allow larger cluster quantities to have more variety in their cluster compositions. But, this also has a drawback in that some models may get biased towards a specific cluster if it appears multiple times in the cluster set and may not best represent data sampling variability. Another option is to create more clusters such that each cluster set is more likely to be a unique combination of clusters and the probability of a single cluster appearing multiple times is diminished. However, we are limited in decreasing the cluster sizes because as the total number of clusters approaches the number of sources, the issue of data leakage resurfaces. Determining the best solution to this issue is a fruitful topic for future research.
There are also avenues for improvement for the statistical framework. Since many of the models share similar characteristics, for example Free or Frozen weights, masking or no masking, etc., we could improve at consolidating this information in the model. Bayesian hierarchical modeling, while more complicated, would allow for flexibility of this nature and could simultaneously address the negative data variance issue through priors.

### v2

@@ 4 Methods - Model Architecture & 3 Training Approaches
We train all models with the deep learning Python package, PyTorch [41] using a sufficient number of epochs to allow full convergence. The Standard models reached convergence within 50 epochs and the TL models reached convergence within 12 epochs. We selected the model instance for evaluation that exhibited lowest loss on the validation data within a cosine annealing cycle. We used cross entropy loss, the ADAM optimizer [42], and a constant batch size of 128 due to its superior performance in preliminary testing.
@@ 5.1 Metrics
As our main goal is to demonstrate the effectiveness and versatility of our evaluation framework, we limit the scope of evaluation of the pickers by displaying just a couple classification metrics and one functional based regression metric. In this section, we briefly outline how we calculate these metrics. To motivate the specific metrics we consider, we note that it is important in evaluation of pickers to consider both the model’s strength in identifying if a P-arrival occurs (e.g., alerting) and, if so, the associated accuracy of the pick (e.g., how close is the predicted pick to the labeled P-arrival).
In general, we implement techniques similar to those found in [1, 7, 32] to obtain model outputs and picks for each test waveform. More specifically, we feed test waveforms through the model in 30 second intervals with 28 seconds of overlap and no end-padding. Then, we take the mean of the model outputs at each time-point to form a single vector of P-arrival probabilities as a function of time for each waveform. Given a defined threshold (described later), we define a pick as the discrete locations in time that a P-arrival is predicted by the model by taking the argmax of groups of samples that contiguously exceed the threshold.
Evaluating the model’s alerting capability is typically done using classification metrics defined from the picks [1, 7, 32, 43]. We define a pick as a true positive ($TP$) if it falls within a true positive window of ±$\pm 0.3$ seconds of the labeled P-arrival. False positives ($FP$) are assigned to picks outside the window or to extra picks that are not true positives. False negatives ($FN$) occur if no picks occur within the window. True negatives are not informative in this context and are not recorded. An example of a model prediction featuring classified picks is shown in Figure 4.
Figure 4: An example of a model output for a test waveform showing the predicted probability of the P-arrival at each time point. The labeled P-arrival is displayed on the x-axis as a blue triangle along with 3 predicted picks (green stars) and labels classifying the pick as either a FP or TP. An inset plot zooms in on the true positive window (orange box: ±$\pm 0.3$ seconds around the labeled P-arrival) to show the TP. Here, an arbitrary threshold of 0.5 (orange line) was used for example purposes.
In our display, we consider summary metrics including $Recall=\frac{TP}{TP+FN}$ and $F1=\frac{2TP}{2TP+FP+FN}$ which are only applied to earthquake waveforms. For noise waveforms, we calculate noise percent correct as the number of noise waveforms containing no predicted picks ($TN_{noise}$) divided by the total number of noise waveforms ($n_{noise}$). That is, $Noise\,\%\,Correct=\frac{TN_{noise}}{n_{noise}}$. We recognize that classifying noise waveforms is a relatively trivial task and that analysts may be interested in different metrics beyond the ones analyzed here.
@@
We define a different threshold for each model instance by choosing the threshold that maximizes the mean of $F1$ and $Noise\,\%\,Correct$ on the validation set in order to seek a balanced performance between noise and earthquake classes. However, this is an arbitrary decision and the choice is typically dependent on the analyst’s goals.
@@ 5.2 Statistical Framework for Evaluation
In this section, we present a statistical framework for comparing models in terms of their performance and uncertainty arising from data sampling and model training. The statistical framework can be thought of as a mixed effects model or a completely randomized design with sub-sampling. It allows us to distinguish performance effects with uncertainty between the 3 models and different data training budgets. Furthermore, we can segregate and estimate the two sources of variation with uncertainty.
@@
where $m=\{1\ldots 3\}$ indexes the 3 model training approaches, $a=\{1,\ldots,5\}$ indexes the 5 quantities of clusters (representing the amount of data) at levels $\{1,3,6,9,12\}$, $d=\{1,\ldots,12\}$ indexes the randomly selected cluster set (i.e., the data choice selected to train the model), and $i=\{1,2,3,4\}$ indexes the 4 model initializations. Here, $\gamma$ is the grand mean, $\mu_{m}$ is the effect of model $m$, $\alpha_{a}$ is the effect of the $a$’th quantity of clusters, and $\theta_{ma}$ is the interaction term.
The uncertainty associated with data sampling and training is segregated into two independent random terms each assuming a zero mean independent and identically distributed Gaussian with variance inherent to a given model and quantity of clusters: $\epsilon^{data}_{mad}\sim\mathcal{N}(0,\sigma^{2data}_{ma})$ and $\epsilon^{train}_{madi}\sim\mathcal{N}(0,\sigma^{2train}_{ma})$. The distribution of the error terms implies a distribution of the metric as $y_{madi}\sim\mathcal{N}(\gamma+\mu_{m}+\alpha_{a}+\theta_{ma},\sigma^{2data}_{ma}+\sigma^{2train}_{ma})$. We recognize that metrics such as recall are bounded between $[0,1]$, so the Gaussian assumption theoretically cannot hold. However, in practice the metrics are calculated from a large number of test waveforms which provides theoretical support for the Gaussian assumption through the central limit theorem. In addition, we find that the variances of these metrics are small compared to the range of the metrics’ bounds. Furthermore, these assumptions are reasonable because the Gaussian distribution is capable of approximating confidence even when the error distribution is not strictly Gaussian. In addition, we verified the assumptions through visual inspection of quantile-quantile plots that compare the sorted residuals as a function of the theoretical quantiles for all metrics of interest. An example of this check for F1 score is shown in Figure 5 demonstrating strong support for the theoretical assumptions.
Figure 5: The Gaussian assumptions are visually verified through quantile-quantile plots. As an example, the sorted training (top) and data (bottom) residuals for F1 score are seen to align closely with the theoretical quantiles as defined by the statistical framework as seen in how the points follow the red line.
Using these models, we estimate the effect of each model and quantity of clusters on the metric by computing the mean over the 12 cluster sets and the 4 initializations. Then, we compute $90\%$ confidence intervals around those estimates using the Gaussian assumptions. Note that the intervals are influenced more heavily by the cluster set (data choice) uncertainty because at each cluster set, we average over the initializations thereby reducing training uncertainty influence.
@@ 6 Results
After training the 720 model instances, computing the performance metrics, and organizing them through the statistical framework, we are ready to compare the performance of the models.
@@
The classification based summary metrics of recall and noise percent correct are presented in Figure 6 along with 90% confidence intervals for the Standard, TL Free, and TL Frozen models showing their performance as a function of quantity of clusters used in training. All models improve in recall as more data is added with largest gains between 1 and 3 clusters followed by diminishing returns. The trend is less consistent for noise percent correct where, in contrast to expectations, the TL Frozen model performs slightly worse with increasing data. The uncertainty in performances decreases for all models as more training data is added.
@@
Figure 6: The statistical framework is used to summarize classification results including earthquake recall (left) and noise percent correct (right) as a function of the quantity of clusters used for training for the 3 training approaches. Vertical bars represent 90% confidence intervals. Largest performance gains occur as the quantity of clusters increases from 1 to 3 and crossing lines indicate different training approaches are superior for different budgets of training data.
@@
The functional regression based metric of cumulative RMSR demonstrating the evaluation’s framework versatility and importance is shown in Figure 7 for the 3 models with 90% confidence intervals (ribbons) for 1, 6, and 12 quantities of clusters. To compute these, we apply the statistical framework (1) treating each discretized element along the domain as a standalone metric. Stark performance differences can be seen between each model especially as more training data is added as seen through tightening confidence intervals. Note that cumulative RMSR ends at $0.3$ seconds because that is the half-width of the true positive window.
@@
In both Figures 6 and 7, we see that there is no clear winner in terms of learning efficiency for all training budgets. The TL Free model demonstrates superiority in the mean performance for recall and cumulative RMSR for the smallest data size of 1 cluster, but is surpassed by the Standard model at 3 clusters and beyond. These results would help practitioners choose the model for their given problem. Depending on the availability of training data and how they prioritize performance metrics, their choice would change.
@@
Figure 7: The statistical framework is used to summarize the regression-based functional metric of cumulative RMSR for 1, 6, and 12 clusters (top, middle, and bottom respectively) for the 3 models. Ribbons indicate 90% confidence intervals. Inset plots zoom into the tails to show how the relative positions of the model curves change as more training data is added indicating that no model is the most efficient learner for all training data budgets.
@@
Our evaluation framework allows for estimation of two sources of performance uncertainty including random data sampling and stochastic training processes. As an example, Figure 8 shows these estimates for F1 for the 3 models across the training data budgets with 90% confidence intervals. We see that training and data variance decrease as more training data is added and we see some differentiation in the estimated variances for each model. The data variance (middle panel) shrinks quickly towards zero as more data is added and is often estimated as negative. These negative estimates are not shown because negative variance estimates are theoretically impossible. This is a known challenge in these classical statistical frameworks, but negative values essentially mean the data variance component is very small relative to the training variance.
@@
The right panel in Figure 8 shows the relative contributions of each source of uncertainty through bar plots where the opaque and transparent bars are the contributions of data and training variances, respectively. The total variance for all models trends downward with more data with diminishing returns. The uncertainty is mostly dominated by the stochastic training processes when quantities of clusters is 3 or more. Practitioners can use this information to target specific sources of variation that they suspect might dominate their problem, perhaps by running more initializations or strategically labeling data.
@@
Figure 8: Training variance (left) and data sampling variance (middle) are estimated with 90% confidence intervals (vertical bars) for F1 for the 3 training approaches. The right panel shows bar plots constructed as the sum of the estimates of training and data variances. The different transparency levels show the contribution of each component where the training contribution is transparent and the data contribution is opaque. Estimating performance uncertainty and attributing the source is informative to practitioners during model selection for setting performance expectations and risk mitigation.
@@
We provide an example that emphasizes the risk of evaluating models without considering uncertainty in performance. We consider the scenario where a practitioner is deciding among the 3 training approaches to incorporate into their workflow. They are equipped with a single cluster set, say 1 cluster’s worth of data. To make their decision, we imagine that they train one instance of each model using the fixed data. We did exactly this in our training design, but we went further by training each model 4 times under different initializations, not just once. As such, we imagine the practitioner is equally likely to get one of the four recall scores we obtained for each model. Thus, with 3 models and 4 equally likely results for each, there are $644^{3}=$ 64 equally likely possible outcomes that the practitioner could realize. If the practitioner ranks each model based on recall score (e.g., first place, second place, third place), due to stochastic training processes, we might expect these rankings to change if the practitioner ran the experiment a second time. The important question then becomes, what is the probability that each model gets each rank?
@@
Using our 3 models each initialized 4 times for a fixed quantity of clusters, we calculate the exact probabilities that each model is ranked in first, second, or third place. Furthermore, since we have 12 cluster sets for each quantity of clusters, we conduct these calculation for each of the 12 sets and take the mean to mitigate data bias. The results of this simulation are presented in Figure 9, where the probability estimates of ranking each model based on recall in first, second, or third place are shown for 1 and 12 cluster quantities.
@@
Figure 9 shows that, due to training and data uncertainty, there is substantial risk of choosing a model that performs more poorly on average. For example, for 1 cluster (top panel), there is a $2.0\%$ chance that TL Frozen has the best recall score. In light of the recall results in Figure 6 (left panel) for 1 cluster, we know that TL Frozen performs substantially worse than the other two models on average. In our case, based purely on recall, it would be a mistake if the practitioner were to choose this model for their workflow. In a similar light, claiming that the TL Frozen model is the state of the art compared to the other two would be a grave mistake as it could encourage technological development in a dead end. Our framework helps guard against such mistakes.
@@
Figure 9: To demonstrate the importance of considering uncertainty in the evaluation framework, the probability of ranking the 3 training approaches based on recall in first, second, or third place is shown for 1 (top) and 12 (bottom) quantities of clusters. Numbers/bars indicate the probability of ranking the associated model in each place if one were to only train one model instance. Numbers less than 100% indicate that mis-ranking the models is a risk which is detrimental to model selection.
@@ 7 Discussion
We present results that will be increasingly relevant for developers of DLMs for the seismic community. In inspecting the results presented in Figure 6, we see overall poor performance in the TL Frozen model. We hypothesized that freezing half the model at the learned weights from the global dataset would produce the best model for small training data amounts. However, the results negate this hypothesis in our case for datasets comprising 10,000 waveforms (approximately the number of waveforms in a single cluster) or more for the PhaseNet architecture. In a similar vein, we see that models trained from scratch perform better than TL when the data size approaches 30,000 waveforms (3 clusters). This suggests that TL might not always lead to better performance and emphasizes the need to continue comparing TL with traditional training approaches.
@@
Our results show that there is no one model that performs the best across all training data quantities. The optimal model choice may depend on the amount of labeled data available for training, especially in low-data scenarios. Another valuable result for the seismic AI community is that expending effort to acquire larger and larger labeled training datasets may be wasteful. This is exhibited by Figure 6 where we see diminishing improvement as the amount of training data increases. This suggest relatively smaller datasets ($\sim$30,000 or $\sim$60,000 waveforms in our case) may be sufficient for reasonable model performance even when training models from scratch. This is good news because large labeled datasets are expensive and sometimes unattainable. We acknowledge that these conclusions may change depending on how a given analyst or developer chooses their data, model architecture, training process, data augmentations, and metric formulations. Estimating learning efficiency could help understand cost and benefit trade off for obtaining more data in this situation.
We emphasized evaluation of models in terms of their learning efficiency while accounting for training and data uncertainty in the metrics of interest. These aspects are valuable for a more wholistic, robust evaluation, but there are more aspects that could supplement our framework by providing context to the raw performance metrics. Some examples of these aspects include the amount of training/computing resources, the intricacy of data augmentations, robustness to noise, predictive uncertainty coverage, architecture complexity, energy efficiency, and model transparency/interpretability. While these important evaluation aspects are not covered in our exposition, our evaluation framework can assist with fair comparison of them going forward.
In this work, we emphasized the importance of controlling data leakage when constructing training splits/budgets to better measure learning efficiency and represent data uncertainty. We constructed our splits by clustering using source locations as a simple and potentially representative approach to how seismic datasets are often geographically constrained. However, clustering on different features informed by domain experts may provide greater control over data leakage, better representation of data sampling procedures, and can target specific performance goals like performance on out-of-distribution test data.
@@
The optimal number of clusters is something to be considered alongside estimating data variance. We noted that some data variances are estimated as negative because they are small relative to training variance. One way to more precisely estimate data variance is to increase the number of cluster sets (i.e. from 12 randomly chosen cluster sets to 30). However, we would not see much variation in the cluster sets for large quantities of clusters (e.g., 9 or 12) because we do not sample with replacement. Sampling with replacement would allow larger cluster quantities to have more variety in their cluster compositions. But, this also has a drawback in that some models may get biased towards a specific cluster if it appears multiple times in the cluster set and may not best represent data sampling uncertainty. Another option is to create more clusters such that each cluster set is more likely to be a unique combination of clusters and the probability of a single cluster appearing multiple times is diminished. However, we are limited in decreasing the cluster sizes because as the total number of clusters approaches the number of sources, the issue of data leakage resurfaces. Determining the best solution to this issue is a fruitful topic for future research.
In our display, we leverage deep ensembles to obtain training uncertainty for performance for a few reasons. Fundamentally, deep ensembles are easy to understand and implement and they only require extending training time. This is in contrast to other uncertainty methods that attempt to estimate prediction uncertainty or approximate the posterior distribution of a given prediction. These methods include Bayesian neural networks [44] and approximations thereof including dropout [45], Stochastic Weight Averaging - Gaussian (SWAG) [46] and many others [47]. However, these approaches enforce distributional assumptions of the model weights, make strong/arbitrary decisions (e.g, dropout), rarely consider different basins of attraction, and are generally more difficult to interpret. Some methods, however, such as MultiSWAG [48] (see [28] for a seismic based implementation), offer a strong balance between the complex Bayesian assumptions and the simpler empirical ensembles and offer strong promise for estimating performance uncertainty.
There are also avenues for improvement for the statistical framework. A Bayesian framework could assist when comparing models with similar characteristics through hierarchical sharing of information. In addition, it could also address the issue of estimating negative data variances through constrained priors.
Inspired by the remarkable success of foundation models (FMs), a subset of AI (typically DL) models trained on vast amounts of data in a self-supervised fashion that can be adapted for various downstream tasks using relatively smaller labeled datasets [49], in the areas of natural language processing and computer vision, FMs are actively being developed for seismology [50, 51, 52, 53]. As the seismic community prepares to assess the utility of seismic FMs, our work provides some useful directions. Because FMs are pre-trained using a large general dataset and fine-tuned (typically weights for the last couple of layers are set to be updated) using tailored datasets to accomplish downstream tasks, our work highlights the importance of assessing the utility of seismic FMs by comparing them to traditional models. Our results also fortify the importance of learning efficiency as an aspect of future FM evaluation. Since the best model depends on the amount of training data, knowing which model or training approach performs better with less data is valuable information during model selection. And, as our results shows, it may not be a simple one-model-takes-all, but rather a dynamic choice for a given situation.